# footnotetext: Corresponding author

### v1

@@ 1 Introduction
In this work, we present Layout-Your-3D, an efficient text-to-3D generation method for compositional 3D scene synthesis with precise control. Specifically, given a text prompt describing few objects and their spatial relationships, we aim to generate these objects and their natural interactions. We note that this is similar to the setting in (Sun et al., 2024) but differs from  (Zhou et al., 2024b; Li et al., 2024a), which focuses on synthesizing complete indoor/outdoor 3D scenes. Our key idea is to utilize a 2D layout as guidance to produce target reference images for 3D scene generation, which not only allows precise user control but also produces more plausible object interactions. We start by collecting a user-provided or LLM-generated 2D layout and predict a reference image depicting the desired instances and their interactions. Using this reference, we propose a carefully designed process to create coarse 3D scenes with roughly reasonable layouts using reconstruction models (Tang et al., 2024a). While these feed-forward reconstruction models are computationally lightweight, they often yield inferior 3D geometry and appearance. To address these limitations, we introduce a two-step disentangled refinement stage to enhance the quality of 3D instances and global layout. First, we develop collision-aware layout optimization to adjust the layouts of the objects to create a globally coherent and visually appealing composition. By incorporating collision awareness, we prevent objects from severe intersections, resulting in a realistic compositional layout. Second, we carry out instance-wise refinement to further improve the geometry and texture of individual objects, which also allows the customization of different instances as an added benefit.
@@ 2 Related Work
Compositional 2D and 3D generation. Generating or reconstructing scenes with multiple objects presents significant challenges in 2D and 3D contexts. In the 2D domain, numerous models (Li et al., 2023; Xie et al., 2023; Zhou et al., 2024a) use 2D bounding box layouts as conditions to guide the generation process. These works perform well in controlling the number of instances and their attributes. Due to the additional depth dimension, generating compositional assets in the 3D domain faces even more significant challenges. Several image-to-3D reconstruction methods, such as that by (Han et al., 2024; Chen et al., 2024b), reconstruct each object in the scene independently and subsequently optimize the global layouts by aligning them with reference images or utilizing additional guidance from diffusion models. Another line of work focuses on compositional text-to-3D generation. For example, COMOGen (Sun et al., 2024) adopts 2D layouts to guide the SDS process to provide accurate control over the generation process. GroundedDreamer (Li et al., 2024b) integrates the core idea of Attend-and-Excite (Chefer et al., 2023) into MVDream (Shi et al., 2024) to facilitate the generation of compositional multi-view images from complex text prompts. Moreover, text-to-3D scene generation methods (Zhou et al., 2024b; Li et al., 2024a) that rely on the score distillation technique often utilize 3D layouts provided by users (Bai et al., 2024) or generated from LLMs to provide coarse compositional spatial information to construct the scene. However, these methods often require hours for generation and may produce unreasonable object interactions or unrealistic appearances. In this work, we leverage efficient reconstruction models to achieve high quality and efficiency, significantly advancing high-fidelity compositional 3D scene generation.
@@ 3.2.2 Instance-wise Refinement
Interactive Editing. Besides 3D scene generation, users can customize the per-instance text prompt $\hat{y}_{i}$ in the instance-wise refinement step for interactive editing. Specifically, we first generate a 3D scene with our coarse 3D generation stage and then stylize each 3D instance in the refinement step, with users’ custom text prompts $\hat{y}_{i}$. This editing application can open up new possibilities for generating extra-complex compositional 3D scenes, with each instance having its attribute and style.
@@ 4.2 Main Results
Comparison methods. We compare Layout-Your-3D with several text-to-3D generation methods, i.e., DreamFusion (Poole et al., 2022), ProlificDreamer (Wang et al., 2023), GaussianDreamer (Yi et al., 2024) and LucidDreamer (Liang et al., 2024). Since our method utilizes an additional 2D layout as a condition, we provide results by both the user-given and LLM-grounded 2D layout. Moreover, we present a comparison of single 3D object generation in Fig. 6.
@@
LucidDreamer 1 hour 22.36% 34.11% 34.84%
@@ A.1 Limitations & Future work
Since Layout-Your-3D relies on reference images to generate compositional 3D assets, it would need deliberately designed 2D layouts to generate extra-complex 3D scene, e.g., having more than 10 instances in the asset. What’s more, as we observed, the SSDS loss, proposed in ComboVerse, is not capable of harmonizing interactions with too many instances involved in.
Thus, in our future work, we will work on deliberately designed modules to achieve more mature text-to-3d layout refinement and generation, without constraints on number of instances or interactions. Also, the application in scene generation of our method is a promising direction, we can extend our method in order to improve the efficiency and controllability of Text-to-3D scene generation.
@@ A.4 Discussion on concurrent methods
GroundedDreamer’s core concept is to marry Attend-and-Excite with MVDream (Shi et al., 2024) to generate compositional and consistent multi-view images, then lift them to 3D assets through an optimization-based method. Though it shows some surprising results and relatively high quality, but due to the limited compositional 3D data used in the training process, the limitations of MVDream still exists, i.e., lacking in prior knowledge for the generation of multiple instances. Also, the pure optimization-based paradigm makes GroundedDreamer low in efficiency, taking nearly 2 hours to generate one single asset. In comparison, Layout-Your-3D needs only about ten minutes (short refinement) or about one hour (extended refinement) to generate one asset. If given more GPU cards, the time cost could be reduced to several minutes and half an hour. Furthermore, since Layout-Your-3D primarily generates individual 3D instances and subsequently combines them, which mitigates the domain gap during the generation process, it effectively helps addressing compositional 3D generation task at its source.
@@
COMOGen takes the same 2D layout setting as adopted in Layout-Your-3D to indicate ideal regions of different instances in the text prompts. However, different from Layout-Your-3D, COMOGen first adopted a Layout-SDS loss to distill the spatial layout knowledge, then uses a Multi-view Control module to ensure multi-view consistency and alleviate Janus Problem, finally proposed a 3D content enhancement module to generate realistic and diverse 3D samples. The Layout-SDS loss could provide a rough estimation of the ideal 3D content, in the following Multi-view Control module, COMOGen adopted a Zero-123 (Liu et al., 2023) multi-view diffusion model to ensure the geometry consistency. But as we mentioned above, current multi-view diffusion models are still subject to the domain gap, thus making it hard to generalize to compositional scenarios, damaging the effectiveness of COMOGen. Moreover, though not presented, the time cost for COMOGen might not be comparable to our Layout-Your-3D, which only costs about 10 minutes to generate a reasonable and plausible compositional 3D scene, as displayed in Fig. 13. Additionally, COMOGen did not show detailed and high quality single object generation results, which constrains the application scenarios of it.
@@
Another line of works, i.e., GALA3D and DreamScene also have similar setting as our Layout-Your-3D. Nonetheless, these works have their own disadvantages. DreamScene is essentially composing different objects into one predefined scene with no specific interaction modeling process during generation, thus will make the resulting scene less realistic and consistent. GALA3D addresses this issue by incorporating 3D layout-conditioned diffusion models as an implicit control over the interactions between objects. However, the produced 3D scene has no background and the quality of objects are relatively inferior, though processed through tedious optimization process. Since the input settings and generated 3D assets are not strictly aligned with ours, we do not make direct comparisons between these methods and our Layout-Your-3D.
@@ A.5 Other reconstruction models
In the coarse 3D generation stage, we could exchange for any 3DGS-based Image-to-3D model for the reconstruction of each instance, e.g., GRM (Xu et al., 2024), Point-to-Gaussian (Lu et al., 2024). What’s more, it is also applicable to replace 3DGS with other 3D representations, like mesh or tetrahedron for advanced or specific usage. Though it will be slightly different to optimize coarse 3D instances in other representation, there are substitutions for the refinement strategies in our pipeline. Also, our final 3DGS-based asset could be converted into mesh with the efficient mesh extraction method proposed in DreamGaussian (Tang et al., 2024b) or advanced mesh extraction methods like SuGaR (Guédon & Lepetit, 2024).
@@ B.1 Validation Set
In accordance with the requirements for compositional 3D generation, each text prompt contains 2 to 5 instances in the compositional 3D asset. This intentional design of our $Comp20$ validation set allows for a comprehensive evaluation on text-to-3D generation methods. The complete prompt list, together with the corresponding 2D layouts (including both user-provided and LLM-generated layouts), are presented in Tab. C.3. Note that in the default experiments we generate reference images with a 512 * 512 resolution, and the 2D layouts are presented in the form of $[x_{1},y_{1},x_{2},y_{2}]$, with each element has a value from range 0 to 512.
@@ B.3 Additional Ablation studies
We conduct more ablation studies to validate the effectiveness of our model selection. First, we compare the visual foundation model used when extracting features for coarse rotation estimation. We choose DINOv2 (Oquab et al., 2023) to estimate the rotation $r_{i}$ since DINOv2 can provide higher-level geometry and semantic information. In comparison, CLIP (Radford et al., 2021) is another model we adopted, but since it is not specifically designed to capture geometry information in images, it causes visible errors in the rotation estimation. Also, we give a detailed time cost of Layout-Your-3D in Tab. 6, including the time cost for each part and different settings, measured on one single RTX6000 GPU.
@@ C.1 Single Object generation
To validate the effectiveness and generalizability of our method, we give more examples in Fig. 16, both the short and extended refinement strategy can bring customized attributes and easily deform the style of 3D assets, thus enabling flexible and controllable generation.
@@ C.3 Compositional generation
(a) The extended refinement strategy.
(b) The short refinement strategy.
Figure 16: Additional visualization results on instance generation with custom text prompts. We show results generated with both the short and extended instance-wise refinement strategies. White images on the top left corner of each figure is the original coarse 3D instances produced by LGM.
@@
tableDetailed prompt list of our proposed $Comp20$ validation set, including overall text prompt $Y_{B}$ and 2D layout set $B$. We provide both the user given 2D layout set and the LLM generated one. Index Text Prompt 1 [A black bottle], with [a white bottle] on one side and [a blue bottle] on the other User [24, 136, 168, 424], [152, 72, 336, 424], [344, 104, 504, 424] LLM [200, 200, 280, 360], [100, 200, 180, 360], [300, 200, 380, 360] 2 [A teddy bear] reading [a book] wearing [blue sunglasses] User [88, 40, 168, 192], [144, 24, 120, 128], [168, 96, 112, 168] LLM [104, 64, 168, 200], [152, 0, 128, 168], [182, 112, 88, 184] 3 [A squirrel] standing on [A box] User [204, 51, 51, 0], [153, 0, 102, 179] LLM [190, 150, 222, 330], [172, 330, 340, 442] 4 [A small blue toy box] riding [a lion] User [0, 153, 102, 0], [102, 204, 153, 153] LLM [0, 140, 80, 10], [120, 200, 168, 168] 5 [A small robot] pushing [a big football] User [51, 153, 179, 102], [194, 102, 204, 128] LLM [75, 250, 225, 200], [225, 275, 200, 200] 6 [A pigeon] having [a bagel] and [a beer] User [102, 102, 358, 358], [230, 307, 358, 384], [358, 204, 435, 384] LLM [40, 88, 232, 416], [240, 360, 368, 416], [376, 168, 488, 416] 7 [An astronaut] sitting on [a sofa] User [153, 102, 384, 409], [51, 153, 460, 384] LLM [165, 130, 347, 392], [80, 100, 432, 372] 8 [A kitten] lying next to [a flower] User [51, 153, 307, 358], [307, 204, 460, 358] LLM [95, 218, 269, 492], [297, 224, 417, 382] 9 [A beagle] wearing [a blue tie] User [117, 51, 394, 460], [230, 204, 281, 435] LLM [88, 96, 424, 464], [208, 288, 312, 464] 10 [A teddy bear] driving [a toy car] User [40, 184, 472, 384], [215, 119, 335, 326] LLM [160, 150, 352, 406], [140, 320, 372, 470] 11 [A lego man] riding [a horse], the horse is wearing [a blue hat] User [ 96, 64, 288, 352], [39, 181, 481, 463], [367, 152, 463, 240] LLM [150, 200, 280, 400], [120, 250, 390, 450], [220, 150, 290, 220] 12 [A blue toy robot] and [a red toy robot] playing [basketball] User [39, 64, 240, 416], [256, 48, 488, 416], [199, 167, 304, 272] LLM [50, 200, 190, 440], [320, 200, 460, 440], [225, 280, 285, 340] 13 [A brown sofa] with [a cushion], [a teddy bear] and [a basketball] on it User [32, 256, 480, 432], [200, 120, 328, 344], [72, 224, 168, 360], [360, 256, 456, 352] LLM [[20, 180, 492, 412], [140, 230, 240, 310], [250, 220, 340, 330], [380, 250, 460, 330] 14 [A dog] taking [a boat] User [24, 256, 488, 384], [160, 104, 336, 304] LLM [186, 150, 326, 290], [50, 250, 462, 780] 15 [A furry fox’s head] with [a cupcake] on it, wearing [red sunglasses] and have [a lego block] on the nose User [40, 64, 480, 512], [168, 72, 336, 216], [64, 256, 440, 376], [168, 264, 256, 407] LLM [176, 128, 336, 288], [224, 88, 288, 152], [200, 160, 312, 200], [240, 208, 272, 240] 16 [A cute stuffed toy crocodile] and [a small stuffed toy rabbit] on [a skateboard] User [40, 360, 472, 432], [72, 72, 296, 376], [296, 200, 424, 376] LLM [40, 304, 480, 432], [80, 112, 280, 344], [296, 176, 440, 336] 17 [A toy dragon] reaching for [a cola] on [a wooden cabinet] User [256, 192, 456, 416], [32, 104, 248, 416], [312, 80, 392, 200] LLM [101, 179, 289, 400], [316, 233, 394, 471], [40, 370, 472, 500] 18 [A tray] on top of [a toy pyramid], on the tray is [a red bottle], [a toy panda] and [an apple] User [96, 208, 424, 471], [56, 56, 136, 184], [408, 120, 480, 184], [16, 192, 504, 240], [192, 16, 336, 192] LLM [150, 300, 350, 500], [140, 260, 360, 320], [160, 100, 220, 260], [240, 130, 320, 260], [330, 120, 390, 180] 19 [A toy heart] and [a white feather] on [a balance scale] User [56, 136, 168, 248], [336, 192, 456, 240], [40, 224, 480, 384] LLM [140, 220, 240, 320], [272, 235, 352, 305], [100, 150, 412, 362] 20 [A green pumpkin] and [a big round orange] under [an umbrella] User [48, 200, 288, 424], [280, 272, 448, 416], [8, 120, 504, 224] LLM [112, 290, 240, 418], [272, 290, 400, 418], [110, 50, 402, 290]

### v2

@@ 1 Introduction
In this work, we present Layout-Your-3D, an efficient text-to-3D generation method for compositional 3D scene synthesis with precise control. Specifically, given a text prompt describing few objects and their spatial relationships, we aim to generate these objects and their natural interactions. We note that this is similar to the setting in (Sun et al., 2024) but differs from  (Zhou et al., 2024b; Li et al., 2024a; Gao et al., 2024), which focuses on synthesizing complete indoor/outdoor 3D scenes. Our key idea is to utilize a 2D layout as guidance to produce target reference images for 3D scene generation, which not only allows precise user control but also produces more plausible object interactions. We start by collecting a user-provided or LLM-generated 2D layout and predict a reference image depicting the desired instances and their interactions. Using this reference, we propose a carefully designed process to create coarse 3D scenes with roughly reasonable layouts using reconstruction models (Tang et al., 2024a). While these feed-forward reconstruction models are computationally lightweight, they often yield inferior 3D geometry and appearance. To address these limitations, we introduce a two-step disentangled refinement stage to enhance the quality of 3D instances and global layout. First, we develop collision-aware layout optimization to adjust the layouts of the objects to create a globally coherent and visually appealing composition. By incorporating collision awareness, we prevent objects from severe intersections, resulting in a realistic compositional layout. Second, we carry out instance-wise refinement to further improve the geometry and texture of individual objects, which also allows the customization of different instances as an added benefit.
@@ 2 Related Work
Compositional 2D and 3D generation. Generating or reconstructing scenes with multiple objects presents significant challenges in 2D and 3D contexts. In the 2D domain, numerous models (Li et al., 2023; Xie et al., 2023; Zhou et al., 2024a) use 2D bounding box layouts as conditions to guide the generation process. These works perform well in controlling the number of instances and their attributes. Due to the additional depth dimension, generating compositional assets in the 3D domain faces even more significant challenges. Several image-to-3D reconstruction methods, such as that by (Han et al., 2024; Chen et al., 2024b), reconstruct each object in the scene independently and subsequently optimize the global layouts by aligning them with reference images or utilizing additional guidance from diffusion models. Another line of work focuses on compositional text-to-3D generation. For example, COMOGen (Sun et al., 2024) adopts 2D layouts to guide the SDS process to provide accurate control over the generation process. GraphDreamer (Gao et al., 2024) establishes a graph to represent objects and their relations. GroundedDreamer (Li et al., 2024b) integrates the core idea of Attend-and-Excite (Chefer et al., 2023) into MVDream (Shi et al., 2024) to facilitate the generation of compositional multi-view images from complex text prompts. Moreover, text-to-3D scene generation methods (Zhou et al., 2024b; Li et al., 2024a; Po & Wetzstein, 2023; Zhang et al., 2024) that rely on the score distillation technique often utilize 3D layouts provided by users (Bai et al., 2024) or generated from LLMs to provide coarse compositional spatial information to construct the scene. However, these methods often require hours for generation and may produce unreasonable object interactions or unrealistic appearances. In this work, we leverage efficient reconstruction models to achieve high quality and efficiency, significantly advancing high-fidelity compositional 3D scene generation.
@@ 3.2.2 Instance-wise Refinement
Interactive Editing. Besides 3D scene generation, users can customize the per-instance text prompt $\hat{y}_{i}$ in the instance-wise refinement step for interactive editing. Specifically, we first generate a 3D scene with our coarse 3D generation stage and then stylize each 3D instance in the refinement step, with users’ custom text prompts $\hat{y}_{i}$. This editing application opens up new possibilities for generating extra-complex compositional 3D scenes, with each instance having its attribute and style.
@@ 4.2 Main Results
Comparison methods. We compare Layout-Your-3D with several text-to-3D methods, i.e., DreamFusion (Poole et al., 2022), ProlificDreamer (Wang et al., 2023), GaussianDreamer (Yi et al., 2024), LucidDreamer (Liang et al., 2024) and GraphDreamer (Gao et al., 2024). Since our method utilizes an additional 2D layout as a condition, we provide results by both the user-given and LLM-grounded 2D layout. Moreover, we present a comparison of single 3D object generation in Fig. 6. Note that we do not compare with works that haven’t released complete or official implementations (e.g., Locally Conditioned Diffusion (Po & Wetzstein, 2023), CG3D (Vilesov et al., 2023)).
@@
LucidDreamer 35 mins 22.36% 34.11% 34.84%
GraphDreamer 2 hours 24.78% 31.18% 44.28%
@@ A.1 Limitations & Future work
Since Layout-Your-3D relies on reference images to generate compositional 3D assets, it would need deliberately designed 2D layouts to generate extra-complex 3D scene, e.g., having more than 10 instances in the asset. What’s more, as we observed, the SSDS loss, proposed in ComboVerse, is not capable of harmonizing interactions with too many instances involved.
Thus, in our future work, we will work on deliberately designed modules to achieve more mature text-to-3d layout refinement and generation, without constraints on the number of instances or interactions. Also, the application in scene generation of our method is a promising direction, we can extend our method in order to improve the efficiency and controllability of Text-to-3D scene generation.
@@ A.4 Discussion on concurrent methods
GroundedDreamer’s core concept is to marry Attend-and-Excite with MVDream (Shi et al., 2024) to generate compositional and consistent multi-view images, then lift them to 3D assets through an optimization-based method. Though it shows some surprising results and relatively high quality, due to the limited compositional 3D data used in the training process, the limitations of MVDream still exist, i.e., lack of prior knowledge for generating multiple instances. Also, the pure optimization-based paradigm makes GroundedDreamer low in efficiency, taking nearly 2 hours to generate one single asset. In comparison, Layout-Your-3D needs only about ten minutes (short refinement) or about one hour (extended refinement) to generate one asset. If given more GPU cards, the time cost could be reduced to several minutes and half an hour. Furthermore, since Layout-Your-3D primarily generates individual 3D instances and subsequently combines them, which mitigates the domain gap during the generation process, it effectively helps address compositional 3D generation task at its source.
@@
COMOGen takes the same 2D layout setting as adopted in Layout-Your-3D to indicate ideal regions of different instances in the text prompts. However, different from Layout-Your-3D, COMOGen first adopted a Layout-SDS loss to distill the spatial layout knowledge, then uses a Multi-view Control module to ensure multi-view consistency and alleviate Janus Problem, finally proposed a 3D content enhancement module to generate realistic and diverse 3D samples. The Layout-SDS loss could provide a rough estimation of the ideal 3D content, in the following Multi-view Control module, COMOGen adopted a Zero-123 (Liu et al., 2023) multi-view diffusion model to ensure the geometry consistency. But as we mentioned above, current multi-view diffusion models are still subject to the domain gap, thus making it hard to generalize to compositional scenarios, damaging the effectiveness of COMOGen. Moreover, though not presented, the time cost for COMOGen might not be comparable to our Layout-Your-3D, which only costs about 10 minutes to generate a reasonable and plausible compositional 3D scene, as displayed in Fig. 13. Additionally, COMOGen did not show detailed and high-quality single object generation results, which constrains the application scenarios of it.
@@
Another line of works, i.e., GALA3D and DreamScene also have similar settings as our Layout-Your-3D. Nonetheless, these works have their own disadvantages. DreamScene is essentially composing different objects into one predefined scene with no specific interaction modeling process during generation, thus making the resulting scene less realistic and consistent. GALA3D addresses this issue by incorporating 3D layout-conditioned diffusion models as an implicit control over the interactions between objects. However, the produced 3D scene has no background and the quality of objects is relatively inferior, though processed through a tedious optimization process. Since the input settings and generated 3D assets are not strictly aligned with ours, we do not make direct comparisons between these methods and our  Layout-Your-3D.
@@ A.5 Other reconstruction models
In the coarse 3D generation stage, we could exchange for any 3DGS-based Image-to-3D model for the reconstruction of each instance, e.g., GRM (Xu et al., 2024), Point-to-Gaussian (Lu et al., 2024). What’s more, it is also applicable to replace 3DGS with other 3D representations, like mesh or tetrahedron for advanced or specific usage. Though it will be slightly different to optimize coarse 3D instances in other representations, there are substitutions for the refinement strategies in our pipeline. Also, our final 3DGS-based asset could be converted into mesh with the efficient mesh extraction method proposed in DreamGaussian (Tang et al., 2024b) or advanced mesh extraction methods like SuGaR (Guédon & Lepetit, 2024).
@@ B.1 Validation Set
In accordance with the requirements for compositional 3D generation, each text prompt contains 2 to 5 instances in the compositional 3D asset. This intentional design of our $Comp20$ validation set allows for a comprehensive evaluation on text-to-3D generation methods. The complete prompt list, together with the corresponding 2D layouts (including both user-provided and LLM-generated layouts), are presented in Tab. D.3. Note that in the default experiments we generate reference images with a 512 * 512 resolution, and the 2D layouts are presented in the form of $[x_{1},y_{1},x_{2},y_{2}]$, with each element has a value from range 0 to 512.
@@ B.3 Additional Ablation studies
We conduct more ablation studies to validate the effectiveness of our model selection. First, we compare the visual foundation model used when extracting features for coarse rotation estimation. We choose DINOv2 (Oquab et al., 2023) to estimate the rotation $r_{i}$ since DINOv2 can provide higher-level geometry and semantic information. In comparison, CLIP (Radford et al., 2021) is another model we adopted, but since it is not specifically designed to capture geometry information in images, it causes visible errors in the rotation estimation.
Also, we give a detailed time cost of  Layout-Your-3D in Tab. 6, including the time cost for each part and different settings, measured on one single RTX6000 GPU.
@@ C.1 Single Object generation
(a) The extended refinement strategy.
(b) The short refinement strategy.
Figure 16: Additional visualization results on instance generation with custom text prompts. We show results generated with both the short and extended instance-wise refinement strategies. White images on the top left corner of each figure are the original coarse 3D instances produced by LGM.
@@
To validate the effectiveness and generalizability of our method, we give more examples in Fig. 16, both the short and extended refinement strategies can bring customized attributes and easily deform the style of 3D assets, thus enabling flexible and controllable generation.
@@ C.3 Compositional generation
Appendix D Additional results & comparisons
D.1 Layout initialization
In this section, we provide more visualization results on the effectiveness of our 3D layout initialization. We compare different 3D layouts generated with LLM inference (Gao et al., 2024; Zhou et al., 2024b) and our strategy in Fig. 18. LLM-generated 3D layouts cannot align with the input text prompts well, thus posing significant optimization problems in the following refinement process. In comparison, our 3D layout initialization strategy benefits from deployed 2D models, and can generate both precise and reasonable layouts. These comparisons reflect the motivation of using 2D blueprint for 3D layout initialization, instead of generating directly from LLMs.
Figure 18: Visualization results of our initialized 3D layout, together with a comparison with LLM-generated coarse 3D layout boxes.
D.2 Comparison with other works
Since ComboVerse has not released the official implementations, we compare  Layout-Your-3Dwith the unofficial version in our experiments. For the unofficial implementation, we use the same initialization strategy as our Coarse 3D Generation Stage, instead of the one presented in ComboVerse. The comparisons are shown in Fig. 19. The left images for each prompt are 3D scenes generated by our implemented ComboVerse and the right ones are results generated with Layout-Your-3D.
Moreover, we compare the results generated with the same prompt by GALA3D and our  Layout-Your-3D in Fig. 20. The figure shows that GALA3D tends to produce unreasonable 3D scenes when given poorly initialized 3D layout boxes. It is worth mentioning that GALA3D takes about 3 hours to generate one scene with 3 instances, which can be inefficient compared to our  Layout-Your-3D  (12 minutes).
Figure 19: Comparison with ComboVerse.
Figure 20: Comparison with GALA3D.
D.3 ablation on rotation estimation
We conduct ablation studies on the effect of visual foundation models used when extracting features for coarse rotation estimation in Tab. 7. The ’Average error’ term in the table represents the average deviation degrees for all instances in the validation set, we calculate and estimate the error in degrees manually. As shown in the table, the deployed visual foundation models significantly reduce rotation error, compared to scenarios where rotation estimation is not applied, thus validating the effectiveness of our coarse rotation estimation strategy.
Table 7: Ablation study on the foundation models used when extracting features for coarse rotation estimation.
Method No estimation CLIP (Radford et al., 2021) DINOv2 (Oquab et al., 2023)
Average error 18.9 10.8 5.5
@@ D.2 Comparison with other works
tableDetailed prompt list of our proposed $Comp20$ validation set, including overall text prompt $Y_{B}$ and 2D layout set $B$. We provide both the user-given 2D layout set and the LLM-generated one. Index Text Prompt 1 [A black bottle], with [a white bottle] on one side and [a blue bottle] on the other User [24, 136, 168, 424], [152, 72, 336, 424], [344, 104, 504, 424] LLM [200, 200, 280, 360], [100, 200, 180, 360], [300, 200, 380, 360] 2 [A teddy bear] reading [a book] wearing [blue sunglasses] User [88, 40, 168, 192], [144, 24, 120, 128], [168, 96, 112, 168] LLM [104, 64, 168, 200], [152, 0, 128, 168], [182, 112, 88, 184] 3 [A squirrel] standing on [A box] User [204, 51, 51, 0], [153, 0, 102, 179] LLM [190, 150, 222, 330], [172, 330, 340, 442] 4 [A small blue toy box] riding [a lion] User [0, 153, 102, 0], [102, 204, 153, 153] LLM [0, 140, 80, 10], [120, 200, 168, 168] 5 [A small robot] pushing [a big football] User [51, 153, 179, 102], [194, 102, 204, 128] LLM [75, 250, 225, 200], [225, 275, 200, 200] 6 [A pigeon] having [a bagel] and [a beer] User [102, 102, 358, 358], [230, 307, 358, 384], [358, 204, 435, 384] LLM [40, 88, 232, 416], [240, 360, 368, 416], [376, 168, 488, 416] 7 [An astronaut] sitting on [a sofa] User [153, 102, 384, 409], [51, 153, 460, 384] LLM [165, 130, 347, 392], [80, 100, 432, 372] 8 [A kitten] lying next to [a flower] User [51, 153, 307, 358], [307, 204, 460, 358] LLM [95, 218, 269, 492], [297, 224, 417, 382] 9 [A beagle] wearing [a blue tie] User [117, 51, 394, 460], [230, 204, 281, 435] LLM [88, 96, 424, 464], [208, 288, 312, 464] 10 [A teddy bear] driving [a toy car] User [40, 184, 472, 384], [215, 119, 335, 326] LLM [160, 150, 352, 406], [140, 320, 372, 470] 11 [A lego man] riding [a horse], the horse is wearing [a blue hat] User [ 96, 64, 288, 352], [39, 181, 481, 463], [367, 152, 463, 240] LLM [150, 200, 280, 400], [120, 250, 390, 450], [220, 150, 290, 220] 12 [A blue toy robot] and [a red toy robot] playing [basketball] User [39, 64, 240, 416], [256, 48, 488, 416], [199, 167, 304, 272] LLM [50, 200, 190, 440], [320, 200, 460, 440], [225, 280, 285, 340] 13 [A brown sofa] with [a cushion], [a teddy bear] and [a basketball] on it User [32, 256, 480, 432], [200, 120, 328, 344], [72, 224, 168, 360], [360, 256, 456, 352] LLM [[20, 180, 492, 412], [140, 230, 240, 310], [250, 220, 340, 330], [380, 250, 460, 330] 14 [A dog] taking [a boat] User [24, 256, 488, 384], [160, 104, 336, 304] LLM [186, 150, 326, 290], [50, 250, 462, 780] 15 [A furry fox’s head] with [a cupcake] on it, wearing [red sunglasses] and have [a lego block] on the nose User [40, 64, 480, 512], [168, 72, 336, 216], [64, 256, 440, 376], [168, 264, 256, 407] LLM [176, 128, 336, 288], [224, 88, 288, 152], [200, 160, 312, 200], [240, 208, 272, 240] 16 [A cute stuffed toy crocodile] and [a small stuffed toy rabbit] on [a skateboard] User [40, 360, 472, 432], [72, 72, 296, 376], [296, 200, 424, 376] LLM [40, 304, 480, 432], [80, 112, 280, 344], [296, 176, 440, 336] 17 [A toy dragon] reaching for [a cola] on [a wooden cabinet] User [256, 192, 456, 416], [32, 104, 248, 416], [312, 80, 392, 200] LLM [101, 179, 289, 400], [316, 233, 394, 471], [40, 370, 472, 500] 18 [A tray] on top of [a toy pyramid], on the tray is [a red bottle], [a toy panda] and [an apple] User [96, 208, 424, 471], [56, 56, 136, 184], [408, 120, 480, 184], [16, 192, 504, 240], [192, 16, 336, 192] LLM [150, 300, 350, 500], [140, 260, 360, 320], [160, 100, 220, 260], [240, 130, 320, 260], [330, 120, 390, 180] 19 [A toy heart] and [a white feather] on [a balance scale] User [56, 136, 168, 248], [336, 192, 456, 240], [40, 224, 480, 384] LLM [140, 220, 240, 320], [272, 235, 352, 305], [100, 150, 412, 362] 20 [A green pumpkin] and [a big round orange] under [an umbrella] User [48, 200, 288, 424], [280, 272, 448, 416], [8, 120, 504, 224] LLM [112, 290, 240, 418], [272, 290, 400, 418], [110, 50, 402, 290]